# MERA: A Comprehensive LLM Evaluation in Russian

### v2

@@ Abstract
Over the past few years, one of the most notable advancements in AI research has been in foundation models (FMs), headlined by the rise of language models (LMs). As the models’ size increases, LMs demonstrate enhancements in measurable aspects and the development of new qualitative features. However, despite researchers’ attention and the rapid growth in LM application, the capabilities, limitations, and associated risks still need to be better understood. To address these issues, we introduce an open Multimodal Evaluation of Russian-language Architectures (MERA), a new instruction benchmark for evaluating foundation models oriented towards the Russian language. The benchmark encompasses 21 evaluation tasks for generative models in 11 skill domains and is designed as a black-box test to ensure the exclusion of data leakage. The paper introduces a methodology to evaluate FMs and LMs in zero- and few-shot fixed instruction settings that can be extended to other modalities. We propose an evaluation methodology, an open-source code base for the MERA assessment, and a leaderboard with a submission system. We evaluate open LMs as baselines and find that they are still far behind the human level. We publicly release MERA to guide forthcoming research, anticipate groundbreaking model features, standardize the evaluation procedure, and address potential societal drawbacks.
@@
Alena Fenogenova1, Artem Chervyakov1,2, Nikita Martynov1, Anastasia Kozlova1, Maria Tikhonova1,2, Albina Akhmetgareeva1, Anton Emelyanov1, Denis Shevelev1, Pavel Lebedev1, Leonid Sinev1, Ulyana Isaeva1, Katerina Kolomeytseva1, Daniil Moskovskiy3, Elizaveta Goncharova2,4, Nikita Savushkin1, Polina Mikhailova1, Denis Dimitrov4,5, Alexander Panchenko3, 4, Sergei Markov1 1SberDevices, 2HSE University, 3Skoltech, 4AIRI, 5SberAI Correspondence: alenush93@gmail.com
@@ 1 Introduction
Figure 1: The MERA benchmark project incorporates 21 tasks within an assessment platform with a fixed experimental pipeline for LLM evaluation for the Russian language.
Recent advancements in Natural Language Processing (NLP) have fostered the development of Large Language Models (LLMs) that exhibit a new level of task-solving performance. The rapid development of their abilities raises questions about their proper evaluation in various aspects, including natural language understanding, ethical judgments, expert knowledge, etc. Also, the latest research Bommasani et al. (2023); Ye et al. (2023) has raised an essential question of standardized evaluation protocol which should cover a broad spectrum of metrics and possible usage scenarios to mitigate the risks associated with AI adoption.
In recent years, AI research has arguably advanced the most in foundation models (FMs) Bommasani et al. (2021), which are models trained on broad data at a scale and adaptable to various downstream tasks. Furthermore, they interact with humans via free-form text instructions, which makes them general-purpose text interfaces for a broad range of scenarios. These models transform how AI systems work and lead to significant societal consequences, which makes it necessary to carefully evaluate their abilities.
@@
The community has addressed the issue with several recently created benchmarks: BIG-bench Srivastava et al. (2023), HELM Bommasani et al. (2023), MT-Bench Zheng et al. (2023) which test models’ expert knowledge, coding skills and advanced abilities lying beyond the scope of classic benchmarks.
@@
However, most of these “new generation” benchmarks are constructed for the English language, while Russian, at this point, lacks a fair instrument for transparent and independent LLM evaluation. Benchmarks like Russian SuperGLUE Shavrina et al. (2020b) and TAPE Taktasheva et al. (2022) are becoming obsolete, and there is a need to cover the evaluation of advanced abilities of new models. Current Russian benchmarks should be revised to satisfy recent trends and challenges and to foster an understanding of LLMs’ behavior.
@@
This paper addresses the problems mentioned above and presents an independent benchmark MERA [^1] https://mera.a-ai.ru/en (Multimodal Evaluation of Russian-language Architectures). This novel benchmark comprises 21 tasks from 11 skill domains in the instruction format, offering a previously unavailable comprehensive standardized evaluation of LLMs and FMs in Russian. The primary objective of this project is to establish a reliable methodology for assessing foundation models in zero-shot and few-shot instruction settings under fixed evaluation scenarios (see Fig. 1 for MERA general idea description). The current benchmark methodology and taxonomy are presented for textual data and its sub-modalities, such as code and formal languages. The methodology is versatile and can be applied to different modalities. We plan to extend benchmark to incorporate other modalities like images and audio in the upcoming MERA releases.
@@
  • we supply a set of baseline solutions, encompassing open source models and human baselines.
@@ 2 Related Work
Benchmarks, such as GLUE Wang et al. (2018) and SuperGLUE Wang et al. (2019), have been the standard evaluation tools for measuring NLP progress for a while. However, recent studies Bender et al. (2021); Elangovan et al. (2021) have criticized their canonical approach for being too shallow and for possible data leakage. In addition, such benchmarks are now considered outdated due to the development of LLMs and FMs for which benchmarks need to be more challenging.
@@
To address these problems, several new benchmarks have been proposed which evaluate LLMs in various settings and scenarios: BIG-bench [^2] https://github.com/google/BIG-bench Srivastava et al. (2023), a massive benchmark comprising more than 200 tasks, is intended to probe LLMs and extrapolate their future capabilities; HELM [^3] https://crfm.stanford.edu/helm/classic/latest/ Bommasani et al. (2023) evaluates LLMs’ generalization abilities in multiple languages and contains an extensive detailed system of metrics for various tasks and scenarios; INSTRUCTEVAL [^4] https://declare-lab.net/instruct-eval/ Chia et al. (2023) provides a comprehensive evaluation methodology for instruction-tuned LLMs. In addition, there is a strong move Hendrycks et al. (2021b); Zhong et al. (2023); Huang et al. (2023) towards assessing a model’s professional knowledge and expertise through exam tasks.
@@
Besides, we see a tendency towards LLM-as-a-judge Zheng et al. (2023) evaluation approach when LLMs (e.g., GPT-4555https://openai.com/research/gpt-4) are used as judges for scoring model answers instead of utilizing automatic metrics (e.g., Accuracy) or human evaluation. However, the standard metrics for generative evaluation were criticized Fomicheva and Specia (2019); Colombo et al. (2022); Chhun et al. (2022); Bommasani et al. (2023) a lot and are not representative. While benchmarks with the systems model-as-a-judge Zheng et al. (2023) [^6] https://lmsys.org could successfully evaluate the model, they have biases, making human judgment more reliable but expensive and unclear in terms of funding.
@@
There have been created benchmarks targeted at even more complex problems such as multimodal knowledge and reasoning Yue et al. (2023), in-context learning Shukor et al. (2023), software development Jimenez et al. (2023), general assistants Mialon et al. (2023); Liu et al. (2023b), social reasoning Gandhi et al. (2023), and alignment skills Ye et al. (2023). An extensive survey of current benchmarks and open challenges is presented in Chang et al. (2023).
@@
However, the limitation of the benchmarks mentioned above is that they are mainly oriented at the English language. As for Russian, there is still a need for a system able to evaluate modern LLM abilities reliably. The main benchmarks for Russian remain Russian SuperGLUE (RSG) Shavrina et al. (2020b), TAPE Taktasheva et al. (2022), and RuCoLA Mikhailov et al. (2022), which do not challenge the modern LLMs enough and do not cover the scope of their recently emerging capabilities. Moreover, following the GLUE and SuperGLUE benchmarks life-cycle, these benchmarks become outdated. More and more tasks in RSG are already solved by the LMs better than by an average human, and only a few remain challenging (e.g., RWSD); the best LMs’ scores on RuCoLA are close to the human results.
As for the modern benchmarks which sufficiently challenge LLMs and FMs’ abilities, there is the rulm-sbs [^7] https://github.com/kuk/rulm-sbs2 benchmark which follows LLM-as-a-judge approach, thus being expensive in evaluation.
To summarize, there is an urgent need for an objective system for independent evaluation of modern LLMs’ abilities in Russian. To address this problem, we propose MERA which brings a platform, a methodology, and 21 tasks for a fair evaluation of foundation models.
@@ 3 Data
The MERA benchmark unites various datasets and benchmarks, which results in 21 tasks covering 11 skill domains for LLM and FM evaluation in Russian.
@@
  • Problem-solving tasks are general intelligence evaluation tasks with a single and non-ambiguous correct solution. They test common intellectual skills (e.g., language understanding, logic, reasoning, and world knowledge) and can be solved by a person without specific training.
@@
Task name Test origin Answer type Skill domain Train Dev Test Prompts
Problem-solving
MathLogicQA
New
Multiple choice
Math, Logic
680 1143 10
MultiQ
TAPE
Free-form
Reasoning
1056 900 5
PARus
RSG
Classification
Common Sense
400 100 500 12
RCB
RSG
Classification
NLI
438 220 438 9
ruModAr
New
Free-form
Math, Logic
6000 6000 5
ruMultiAr
New
Free-form
Math
1039 1024 6
ruOpenBookQA
TAPE
Multiple choice
World Knowledge
2338 400 10
ruTiE
New
Classification
Reasoning,
Dialogue Context,
Memory
@@
ruWorldTree
TAPE
Multiple choice
World Knowledge
115 525 10
RWSD
RSG
Classification
Reasoning
606 204 260 10
SimpleAr
New
Free-form
Math
1000 1000 6
Exam-based
BPS
New
Classification
Code, Math
250 1000 8
CheGeKa
TAPE
Free-form
World Knowledge
29376 416 4
LCS
New
Free-form
Code, Math
320 500 6
ruHumanEval
New
Free-form
Code
164 164 10
ruMMLU
New
Multiple choice
Reasoning
10033 961 5
USE
New
@@
Multiple choice, free-form, matching
Exam
2622 900 900 3x5*
Ethics
ruDetox
Russian dataset
Free-form
Ethics
6948 800 8
ruEthics
TAPE
Classification
Ethics
1935 5x3*
ruHateSpeech
New
Classification
Ethics
265 10
ruHHH
Translated
Classification
Ethics
178 10x3*
@@
Table 1: The MERA tasks outline. The Train, Dev, and Test columns show the sizes of the dataset splits (where “–” means the absence of such split). “Validation” split is an alias for “Dev” one. The column Prompts shows the number of unique instruction prompts for each task (see Sec. 4.1 for the details). * The ruEthics, ruHHH, and USE datasets include several sub-tasks with different sets of prompts for each of them. In these cases, the number of prompts is presented as the number of prompts per sub-tasks multiplied by the number of sub-tasks.
@@
Based on the taxonomy above and modern practices Chia et al. (2023); Srivastava et al. (2023), we chose 21 tasks which cover main intelligence skills and knowledge domains and are complex enough for current LLMs, and can be evaluated via automatic metrics. The tasks are formulated in the instruction format, targeting various answer types: classification problems (8 tasks), multiple choice questions (5 tasks), free-form answer (9 tasks), and matching (1 task).
All tasks comprise at least a test set with closed answers, which is necessary for model evaluation with MERA. The exception is the diagnostic datasets the answers to which are made public since they are not used in the final assessment. For some tasks, we also include public training and validation sets, even though they are not requisite for MERA evaluation. We do this for several reasons: 1) train and development sets can be used as a source for few-shot examples; 2) for general consistency of the sets that were adapted from other publicly available datasets (e.g., RSG, BIG-bench). We invite the community to use these datasets for general research purposes. Nevertheless, in line with the BIG-bench paradigm Srivastava et al. (2023) and according to the rules of our leaderboard, it is strictly prohibited to use these as train data before submission.
@@
Some tasks were created from scratch for MERA, while others represent converted, adapted, and enriched versions of previously published Russian and translated English datasets. It should be noted that we adapted train and validation data for some tasks (e.g., ruMMLU) while creating an entirely new test set to ensure no data leakage.
Tab. 1 gives an outline of general task information, test data origin, answer type, and the dataset split sizes. The detailed description of the tasks and the corresponding datasets can be found in App. A.
@@
We embed all the data into an instruction format using the following JSON structure for each sample:
@@
  • inputs contains the sample information (problem details);
  • outputs (available only for train and dev sets or the diagnostic tasks) contain the golden answer [^8] Except for ruEthics, where “outputs” correspond to five ethical norms.;
  • meta is a dictionary containing the task id and other relevant meta-information.
@@ 4.1 Methodology
The paper introduces a methodology to evaluate FMs and LMs in zero- and few-shot fixed instruction settings that can be extended to other modalities. The benchmark is designed as a black-box test to ensure the exclusion of any potential data leakage from the test set.
The evaluation procedure is designed to match the instruction format of task datasets under zero- and few-shot settings and is based on the lm-harness framework Gao et al. (2022) [^9] https://github.com/EleutherAI/lm-evaluation-harness/tree/v0.3.0.
There are two strategies to assess the performance of language models used in this framework. The first approach takes the continuation of the input string with the largest log-likelihood, where log-likelihood is computed as a sum of per-token log probabilities. The second approach is greedy generation where the generation process continues greedily until the predefined stopping criteria are met. Tab. 2 shows the distribution of the benchmark tasks over the two assessment strategies.
@@
Log-likelihood
MathLogicQA
5 Acc
PARus
0 Acc
RCB
0 Acc / F1 macro
ruOpenBookQA
5 Acc / F1 macro
ruTiE
0 Acc
ruWorldTree
5 Acc / F1 macro
RWSD
0 Acc
BPS
2 Acc
LCS
2 Acc
ruMMLU
5 Acc
ruEthics
0 5 MCC
ruHateSpeech
0 Acc
ruHHH
0 Acc
Greedy generation
MultiQ
0 EM / F1
ruModAr
0 Acc
ruMultiAr
5 Acc
SimpleAr
5 Acc
CheGeKa
4 EM / F1
ruHumanEval
0 Pass@k
USE
0 Grade norm
ruDetox
0 J(STA, SIM, FL)
@@
Performance of LLMs and FMs may deviate substantially depending on the prompt used Radford et al. (2019); Jiang et al. (2020); Shin et al. (2020); Gao et al. (2021); Schick and Schütze (2021); Lu et al. (2022). MERA seeks to evaluate LLMs’ abilities in a fixed experimental setup, mitigating the influence of prompt selection. Employing the methodology proposed in Li et al. (2023), we manually designed a variative set of prompts for each task as follows.
Each dataset is accompanied by a number of its general task formulations (instructions or prompts) that differ in complexity. The exact number of instructions for the task may differ, considering the complexity and diversity of samples in a dataset. It was estimated from an empirical analysis of the tasks through a series of experiments on training or validation sets or examination of their structure. The upper bound for the number of instructions is set to 12. The exact number of instructions is provided in Tab. 1.
Several annotators were involved in manual prompt creation to mitigate bias and ensure impartiality. Instructions are designed universally, without any reference to data or model architecture, and evenly distributed among data samples in a deterministic manner to ensure equal support for each instruction within the task.
We also define the number of shots for each task and fix the choice of the few-shot examples for further reproducibility. See Tab. 1 for the exact few-shot number and App. B for the motivation of the choice. When creating a prompt in a few-shot setting, we use instructions only for the first shot. The remaining $k-1$ shots (where $k$ is the number of few-shot examples) and the test example are formatted automatically in the generic format incorporated in our adaptation of the lm-harness.

### v3

@@ Abstract
Over the past few years, one of the most notable advancements in AI research has been in foundation models (FMs), headlined by the rise of language models (LMs). However, despite researchers’ attention and the rapid growth in LM application, the capabilities, limitations, and associated risks still need to be better understood. To address these issues, we introduce a new instruction benchmark, MERA, oriented towards the FMs’ performance on the Russian language. The benchmark encompasses 21 evaluation tasks for generative models covering 10 skills and is supplied with private answer scoring to prevent data leakage. The paper introduces a methodology to evaluate FMs and LMs in fixed zero- and few-shot instruction settings that can be extended to other modalities. We propose an evaluation methodology, an open-source code base for the MERA assessment, and a leaderboard with a submission system. We evaluate open LMs as baselines and find they are still far behind the human level. We publicly release MERA to guide forthcoming research, anticipate groundbreaking model features, standardize the evaluation procedure, and address potential ethical concerns and drawbacks.
@@
Alena Fenogenova1, Artem Chervyakov1,2, Nikita Martynov1, Anastasia Kozlova1, Maria Tikhonova1,2, Albina Akhmetgareeva1, Anton Emelyanov1, Denis Shevelev1, Pavel Lebedev1, Leonid Sinev1, Ulyana Isaeva1, Katerina Kolomeytseva1, Daniil Moskovskiy3,4, Elizaveta Goncharova2,4, Nikita Savushkin1, Polina Mikhailova1, Anastasia Minaeva1, Denis Dimitrov4, Alexander Panchenko3,4, Sergei Markov1 1SaluteDevices, 2HSE University, 3Center for Artificial Intelligence Technology, 4AIRI Correspondence: alenush93@gmail.com
@@ 1 Introduction
Figure 1: The MERA benchmark illustration. The benchmark incorporates 21 tasks covering 10 skills within an assessment platform with a fixed experimental pipeline for LLM evaluation for the Russian language.
Recent advancements in NLP have led to the emergence of powerful Large Language Models (LLMs), which showcase unprecedented task-solving capabilities. In recent years, AI research has made notable progress in foundation models (FMs) Bommasani et al. (2021) trained on extensive data and adaptable to various downstream tasks. Interacting with humans through free-form text instructions, these models serve as versatile text interfaces for multiple scenarios, transforming the landscape of AI systems. The swift evolution of models provokes critical questions regarding their comprehensive evaluation, spanning natural language understanding, ethical considerations, expert knowledge, etc. The most recent research Bommasani et al. (2023); Ye et al. (2023) underscores the crucial need for a standardized evaluation protocol encompassing diverse metrics and potential usage scenarios to address risks associated with AI adoption.
@@
The community has addressed the issue with several recently created benchmarks: BIG-bench Srivastava et al. (2023), HELM Bommasani et al. (2023), MT-Bench Zheng et al. (2023) which test models’ expert knowledge, coding skills and advanced abilities beyond the scope of classic GLUE-style Wang et al. (2018) benchmarks.
@@
However, most of these recent benchmarks are constructed for the English language. Russian, at this point, lacks a fair instrument for transparent and independent LLM evaluation. Benchmarks like Russian SuperGLUE Shavrina et al. (2020b) and TAPE Taktasheva et al. (2022) do not cover the entire scope of modern LLM abilities. Current Russian benchmarks should be revised to satisfy recent trends and challenges and to foster an understanding of LLMs’ behavior.
@@
This paper addresses the problems above and presents the benchmark MERA [^1] https://mera.a-ai.ru/en. The project, led by AI Alliance Russia [^2] https://a-ai.ru/, represents a pioneering collaboration between industry and academic partners. MERA plays a crucial role in fostering cohesion between the scientific community and industry, thus maintaining the benchmark’s independence and impartiality. This novel benchmark comprises 21 tasks covering 10 skills in the instruction format, offering a comprehensive standardized evaluation of LLMs and FMs in Russian. The primary objective of this project is to establish a reliable methodology for assessing foundation models in zero-shot and few-shot instruction settings under fixed evaluation scenarios (see Fig. 1 for MERA general idea description). The current benchmark methodology and taxonomy are presented for textual data and sub-modalities, such as code and formal languages. The methodology is versatile and can be applied to different modalities. We plan to extend the benchmark to incorporate images and audio in the upcoming MERA releases.
@@
  • we supply a set of baseline solutions, including open-source models and human baselines.
@@ 2 Related Work
Benchmarks, such as GLUE Wang et al. (2018) and SuperGLUE Wang et al. (2019), have been the standard evaluation tools for measuring NLP progress for the last 5 years. However, recent studies Bender et al. (2021); Yu et al. (2023); Arora and Goyal (2023) have criticized their canonical approach for being too shallow and for possible data leakage. Moreover, given the development of LLMs and FMs, current benchmarks are now considered not challenging enough for modern LLMs, which have outperformed the human level for most of the included tasks. Thus, there is a need for more challenging benchmarks that follow the instruction format relevant to modern instruction-based models.
@@
To address these problems, the community has proposed several new benchmarks evaluating LLMs in various settings and scenarios: BIG-bench [^3] https://github.com/google/BIG-bench Srivastava et al. (2023), a massive benchmark comprising more than 200 tasks, is intended to probe LLMs and extrapolate their future capabilities; HELM [^4] https://crfm.stanford.edu/helm/classic/latest Bommasani et al. (2023) tests LLMs’ generalization abilities in multiple languages and contains an extensive detailed system of metrics for various evaluation scenarios; InstructEval [^5] https://declare-lab.github.io/instruct-eval Chia et al. (2023) provides a comprehensive evaluation methodology for instruction-tuned LLMs. In addition, there is a strong move Hendrycks et al. (2021b); Zhong et al. (2023); Huang et al. (2023) towards assessing a model’s professional knowledge and expertise through exam tasks.
@@
Besides, there is a trend Zheng et al. (2023); Kocmi and Federmann (2023a, b) on using the LLM-as-a-judge evaluation approach when LLMs (e.g., GPT-4666https://openai.com/research/gpt-4 Achiam et al. (2024)) are used to score models in a generation setup instead of utilizing automatic metrics (e.g., BLEU Papineni et al. (2002)) or human evaluation. However, the standard metrics for generative evaluation were criticized Fomicheva and Specia (2019); Colombo et al. (2022); Chhun et al. (2022); Bommasani et al. (2023) a lot for being not representative enough. While benchmarks with the systems model-as-a-judge Zheng et al. (2023) [^7] https://lmsys.org could successfully evaluate a model, they have biases, making human judgment, which is expensive and unclear in terms of funding, more reliable.
@@
Several benchmarks were introduced to target at even more complex problems, such as multimodal knowledge and reasoning Yue et al. (2023), in-context learning Shukor et al. (2023), software development Jimenez et al. (2024), general assistants Mialon et al. (2023); Liu et al. (2023b), social reasoning Gandhi et al. (2023), and alignment skills Ye et al. (2023). An extensive survey of current benchmarks and open challenges is presented in Chang et al. (2024).
@@
However, one of the limitations of the benchmarks mentioned above is that they are mainly oriented on the English language. As for Russian, there is still a need for a system able to evaluate modern LLM abilities reliably. The main benchmarks for Russian remain Russian SuperGLUE (RSG) Shavrina et al. (2020b), TAPE Taktasheva et al. (2022), and RuCoLA Mikhailov et al. (2022), which do not challenge the modern LLMs enough or cover the scope of their recently emerging capabilities (e.g., expertise in science fields or coding skills). More and more tasks in RSG are already solved by LMs better than by an average human, and only a few remain challenging (e.g., RWSD); the best LMs’ scores on RuCoLA are close to the human results. As for the modern benchmarks that sufficiently challenge LLMs and FMs’ abilities, there is the rulm-sbs [^8] https://github.com/kuk/rulm-sbs2 benchmark which follows the LLM-as-a-judge approach, thus being expensive in evaluation.
To summarize, there is an urgent need for an objective system to evaluate modern LLMs’ abilities in Russian independently.
@@ 3 Data
The MERA benchmark unites various datasets and benchmarks, which results in 21 tasks covering 10 skills for LLM and FM evaluation in Russian.
@@
  • Problem-solving tasks are general intelligence evaluation tasks with a single and non-ambiguous correct solution. They test common intellectual abilities and can be solved by a person without specific training.
@@
Task name Test origin Answer type Skills Train* Dev Test Prompts
Problem-solving MathLogicQA New Multiple choice Mathematics, Logic 680 1143 10
MultiQ TAPE Free-form Reasoning 1056 900 5
PARus RSG Classification Common Sense 400 100 500 12
RCB RSG Classification NLI 438 220 438 9
ruModAr New Free-form Mathematics, Logic 6000 6000 5
ruMultiAr New Free-form Mathematics 1039 1024 6
ruOpenBookQA TAPE Multiple choice World Knowledge 2338 400 10
ruTiE New Classification Reasoning,
Dialogue System
@@
ruWorldTree TAPE Multiple choice World Knowledge 115 525 10
RWSD RSG Classification Reasoning 606 204 260 10
SimpleAr New Free-form Mathematics 1000 1000 6
Exam-based BPS New Classification Algorithms 250 1000 8
CheGeKa TAPE Free-form World Knowledge 29376 416 4
LCS New Classification Algorithms 320 500 6
ruHumanEval New Free-form Computer Code 164 164 10
ruMMLU New Multiple choice Reasoning 10033 961 5
@@
USE Adapted Multiple choice, free-form, matching Reasoning, NLI, World Knowledge 2622 900 900 3x5**
Ethics ruDetox Adapted Free-form Ethics 6948 800 8
ruEthics TAPE Classification Ethics 645 5x3**
ruHateSpeech New Classification Ethics 265 10
ruHHH Adapted Classification Ethics 178 10x3**
@@
Table 1: The MERA tasks outline. Test origin discloses the source of the dataset test split. The Train, Dev, and Test columns show the sizes of the dataset splits (“–” means the absence of the split). “Validation” split is an alias for “Dev” one. The column Prompts shows the number of unique instruction prompts for each task (see Sec. 4.1 for the details). * For LCS, ruHumanEval, ruMMLU and ruModAr we report the size of the public test split. ** For ruEthics, ruHHH, and USE datasets we report the number of prompts per sub-tasks multiplied by the number of sub-tasks.
@@
Based on the taxonomy above and modern practices Chia et al. (2023); Srivastava et al. (2023), we chose 21 tasks that test advanced LMs and FMs’ capabilities that can be evaluated via automatic metrics, which we attribute to 10 skills derived from categorizations described in Wang et al. (2018); Shavrina et al. (2020b); Srivastava et al. (2023). The tasks are formulated in the instruction format, targeting various answer types: classification problems (9 tasks), multiple choice questions (5 tasks), free-form answers (8 tasks), and matching (1 task). See Tab. 1 for the general task information; the detailed task description can be found in App. A.
All tasks comprise at least a test set with closed answers. The exception is the diagnostic datasets whose answers are made public since they are not used in the final assessment.
For four datasets (LCS, ruHumanEval, ruMMLU, ruModAr), we also release a public test set adapted from the original public tests of the corresponding datasets. We invite the community to use these datasets as public tests for general research purposes.
For some other tasks, we additionally publish sets marked as training and validation (or dev) sets. We do this for the following reasons: 1) these sets can be used as a source for few-shot examples; 2) for the general consistency of the sets adapted from other publicly available datasets (e.g., RSG, BIG-bench).
Nevertheless, in line with the BIG-bench paradigm Srivastava et al. (2023) and according to the rules of the leaderboard, it is prohibited to use benchmark data in model training.
@@
Some tasks were created from scratch for MERA, while others represent adapted and enriched versions of previously published Russian and translated English datasets. For some tasks, we adapted only public test data (e.g., ruMMLU) while creating a new test set to avoid data leakage. It should also be noted that despite using the translated or adapted data, we paid special attention to incorporating culture-specific aspects in the benchmark datasets (see App. B for more details).
@@
We embed all the data into an instruction format using the following JSON structure for each example:
@@
  • inputs contains the sample information (data);
  • outputs (available for the train, dev, and public test sets or the diagnostic tasks) contain the golden answer [^9] Except for ruEthics, where “outputs” correspond to five ethical norms.;
  • meta is a dictionary containing the sample id and other relevant meta-information.
@@ 4.1 Methodology
The paper introduces a methodology for FMs and LMs evaluation in zero- and few-shot fixed instruction settings that can be extended to other modalities. The benchmark is designed as a private test to exclude potential data leakage from the test set.
The evaluation procedure is designed to match the instruction format of task datasets under zero- and few-shot settings and is based on the lm-eval framework [^10] https://github.com/EleutherAI/lm-evaluation-harness/tree/v0.3.0 Gao et al. (2022); Biderman et al. (2024).
There are two strategies to assess the performance of language models used in this framework. The first approach takes the continuation of the input string with the largest log-likelihood, where log-likelihood is computed as a sum of per-token log probabilities of the continuation, as specified in Eq. 1.
$LL(cont)=\sum_{i=|ctx|+1}^{|ctx|+|cont|}\operatorname{log_{p_{\theta}}}(x_{i}|x_{<i})$ (1)
where $|ctx|$ and $|cont|$ are the token lengths of the initial prompt and the continuation, respectively.
The second approach is greedy generation, where the generation process continues greedily until the predefined stopping criterion is met (by default, until the EOS token is generated).
We use the log-likelihood strategy for the classification and multiple-choice tasks where a certain number of classes limits the set of answers as we want to test the model’s actual skills, not its ability to follow the exact task format (spaces, commas, etc.). The generation strategy is used for the rest of the tasks with a more complex answer structure (see Tab. 2 for the specification).
@@
Log-likelihood MathLogicQA 5 Acc
PARus 0 Acc
RCB 0 Acc / F1 macro
ruOpenBookQA 5 Acc / F1 macro
ruTiE 0 Acc
ruWorldTree 5 Acc / F1 macro
RWSD 0 Acc
BPS 2 Acc
LCS 2 Acc
ruMMLU 5 Acc
ruEthics 0 5 MCC
ruHateSpeech 0 Acc
ruHHH 0 Acc
Greedy generation MultiQ 0 EM / F1
ruModAr 0 EM
ruMultiAr 5 EM
SimpleAr 5 EM
CheGeKa 4 EM / F1
ruHumanEval 0 Pass@k
USE 0 Grade norm
ruDetox 0 J(STA, SIM, FL)
@@
Performance of LLMs and FMs may deviate substantially depending on the prompt used Radford et al. (2019); Jiang et al. (2020); Shin et al. (2020); Gao et al. (2021); Schick and Schütze (2021); Lu et al. (2022). MERA seeks to evaluate LLMs’ abilities in a fixed experimental setup. We mitigate the influence of prompt selection by fixing a prompt (or instruction) for each sample and evenly distributing them among data examples (see Sec. 3 for the exact format). The latter is formatted in the instruction format before being passed to the model. Employing the methodology proposed by Li et al. (2023), we manually designed a variation set of prompts of various difficulties for each task. The prompt number for the task depends on the complexity and diversity of samples in a dataset and is provided in Tab. 1. It was experimentally estimated from an empirical task analysis. Several annotators were involved in manual prompt creation to mitigate bias and ensure impartiality. Instructions are designed universally without any reference to data or model architecture.
We also define the number of shots for each task and fix the choice of the few-shot examples for further reproducibility. See Tab. 2 for the exact few-shot number and App. D for the motivation of the choice. When creating a prompt in a few-shot setting, we use instructions only for the first shot. The remaining $k-1$ shots (where $k$ is the number of few-shot examples) and the test example are formatted automatically in the generic format incorporated in our adaptation of the lm-eval.